OpenStudy (weasel123):

why is 2 √x+5+7=11 extraneous? will give medal

zepdrix (zepdrix):

It seems that x=-1 solves this equation,\[\large\rm 2\sqrt{-1+5}+7\quad= 2\sqrt{4}+7\quad=2\cdot2+7\quad=4+7\quad=11\]This solution is not extraneous. So I'm not really sure what you're asking.